STAFF RESOURCES

Minutes of November 10, 2010 meeting

Present:
Edward Suarez, Jamie Maguire, Kimberly Kaiser (Minutes), Alex Arencibia, Bill Lynch, Andy Martinez, Melody Tomaszewicz
Absent:
Theresa Macklin, Robert Krack, Fred Onorato, Shirley Peck, Mary Brodman

Agenda

Information Sharing

Eddie (LSM): Facilities has removed the unused current periodicals shelving from the third floor. This open space will be repurposed as plans for the floor further develop. Eddie continued shifting the third floor stacks and has reached the journal of C titles. Space becomes problematic at "The Journal of 'B'." The Methods of Biochemical Analysis has been integrated into the stacks causing a shift of about 10 sections. LSM plans to shelf read the second floor over intersession. Eddie coordinated with Rob to send 42 shelves to Physics. Melanie at Math is in need of bookends.

Melody (DTS): Alex is relocating microfilm and integrating microfilm transferred from Douglass.

Jamie (DTS): Three different collections of microfilm are being weeded and moved to the Annex. The Chinese and Korean East Asian Library periodicals are being moved downstairs by the bound periodicals. A new sub location is being created for these items because they are shelved by call number. About 100 items are being transferred from Music to Annex. Weeding is continuing for items from 2B, Camden and Newark. Some of these items are also being transferred to Annex.

Bill (Douglass): The shift in the west room is continuing. The space next to the Fordham Lab will be used as a conference room. The students have been trained on shifting so the pace of the shift is now increasing. Bill has been able to keep up with the in-transit lists for Douglass.

Kim (Kilmer): The business transfers from Alexander to Kilmer have been put on hold as DTS is waiting for selector decisions on the material. Space is available at Kilmer when the project resumes. The quarterly missing book reports have been printed and are ready for searching. During the month of October, the amount of shelf reading has increased. Specifically in the stacks, a total of 57 hours of shelf reading was completed, encompassing 30% of that collection.

Andy (Douglass): A weeding project for examining duplicate items has begun in the stacks (call numbers L-Z). Before withdrawal items will be verified to be on the shelf in the given locations and the copies in the best condition will be kept. All withdrawn items will be sent to Better World Books.

Alex (Dana): A mini shift is being done to incorporate the reference overflow. Construction within the library continues and has been noisy at times.

Rob (Alexander) (Information sent by email after meeting) - The government documents sections is now in order again after the carpet installation project. The backlogged microfiche is now shelved. Government Document shelving is now being completed on a weekly basis. Some weeding in the Art library was done with other members of the staff at Alexander pitching into the project.

Shelving Non Stack Materials - Math library follow up

In previous meetings, the question of where to shelve VHS, DVD and other accompanying materials was discussed. Items will be cataloged into collections, like STACKS-COM, only if the library has this collection preexisting. This question specifically pertained to the Math Library. Math has a Stack-Com collection so materials that fit this collection type will be cataloged for this sub-location. Andy suggested it would be better to place these collections in a public area with clear labeling.

Reviewing Print Serial Cancellations

Andy sent out a list of periodicals that will be canceled at each library to the collection management staff. The list is useful because it will help in planning ahead for space that will become available once these titles have been canceled. DTS will receive letters from the publishing company when the subscriptions have ended. The record will be updated and the necessary collection management staff will be notified. The use of Director's Station was looked into in the planning of collection shifts for serials. Upon looking into the issue it should be noted that Collection Management staff should compile a cancelled serials list in Director's Station. Titles that are cancelled in bound form may appear on an active list as well because RUL still has subscription to the title. Director's Station is updated daily and any changes to the title status will be reflected in its reports.

Emergency Water Supplies

Eddie sent out an email asking staff to do an inventory of water emergency supplies. Kilmer is the only location in need of supplies. Plastic bags will be ordered for Kilmer.

Next meeting

Wednesday, December 8, 2010
Time: 2:30-4:30 pm Sakai Chat Room.



 
URL: http://www.libraries.rutgers.edu/rul/staff/groups/collection_mgt_group/minutes/collection_mgt_group_10_11_10.shtml
Website Feedback  |  Privacy Policy

© Copyright 1997-2013, Rutgers University Libraries   (Further Copyright Information)